Qhen finding the measure of the interior angle and finding X how do you find
J,K,L,M If J=(3x-6) K=(x+10) L=2x and M=(2x-8)?
i used the equation-n-2(180) add up all my x's and numbers and got 8x-4=360
+4 +4 and got 8x=364 then divided that by 8 to get X=44.5

Answer by KMST(5396)

You got it right. You could go a little further and calculate the angles if needed.
If JKLM is a polygon, it has 4 sides/4 angles, so n=4, and the sum of its interior angle measures (in degrees) is n-2(180) = 2(180) = 360.
With J=3x-6, K=x+10, L=2x, and M=2x-8, your equation would start as
(3x-6)+(x+10)+2x+(2x-8) = 360
It would simplify into
8x-4 = 360
and the solution would be x=45.5 as you said,
Then, you could calculate the measures of the angles (in degrees):
J = 3x-6 = 3(45.5)-6 = 136.5 - 6 = 130.5 = 130 degrees 30 minutes
K = x+10 = 45.5 + 10 = 55.5 = 55 degrees 30 minutes
L = 2x = 2(45.5) = 91
and M = 2x-8 = 2(45.5) - 8 = 83